Week of 060626

Open Actions from last week:

Chair: James

SMOD: Veronique

Monday:

Log: R-GMA problems over the weekend due to bad configured sites - no data

New Actions:

  • Maite to follow up with R-GMA team on how to renamble things
  • Intervention on c2atalas, c2cms, c2sc4 clusters (Jan)

Discussion:

  • ATLAS have been running at up to 650MB/s

Tuesday:

Log: Castor SRM down for 7 hours.

New Actions:

  • Create alarm to check for request spool filling up (Jan)

Discussion:

  • Castor problem due to lack of inodes on request spool - cronjob was missgin - now in place

Wednesday

Log:

Actions:

  • Many tickets open with no action on RB - Yvan should follow up - and be asdked to come to wednesday 9am meeting to report on progress
  • Infosys mising info/instabilities - james/Jan/Thorsten
Discussion:
  • Review of action Items from GGUS

Thursday

Log:

Actions:

Discussion:

Friday

Log:

Actions:

Discussion:

Edit | Attach | Watch | Print version | History: r5 < r4 < r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r5 - 2007-02-19 - FlaviaDonno
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback